A vessel of 2×10⁻³m³ volume contains 0.036 kg of Argon at 300K. Calculate its pressure using ideal gas equation. ( Atomic mass of Ar = 39.9u).
1
Expert's answer
2020-12-21T03:53:45-0500

V= 2x10-3m3 = 2dm3

m= 0.036Kg = 36g

T= 300K

R= 0.0821L.atm/k.mol

number of moles, n= mass/Molar mass = 36/39.9= 0.9mol

PV= nRT

P= nRT/V

p= 0.9x0.0821x300/2

P= 11.08atm